# Morphing Tokens Draw Strong Masked Image Models

### v2

@@ 5.3 ADE20K Semantic Segmentation
5.4 Transfer learning
5.4.1 iNaturalist datasets.
We verify the improved transferability of our pre-trained model. We compare fine-tuning accuracies of the baseline and our proposed model on iNaturalist 2018, iNaturalist 2019, and mini iNaturalist 2021 [45], which are highly imbalanced with different numbers of images per class. All the models are ViT-B/16 with a resolution of 224224. Following the protocol [25], we perform grid searches on learning rates and weight decay and report the maximum accuracy and the mean and standard deviation of the accuracies. Table 8 shows our DTM loss significantly improves the baseline with large margins, which reveals enhanced transferability.
5.4.2 Fine-Grained Visual Classification (FGVC) datasets.
We further validate the transferability of our method. Following the evaluation protocol as above, we conduct comparisons on FGVC datasets. Specifically, we evaluate fine-tuning accuracies on Birds, CUB-200, CIFAR-10, CIFAR-100, and Dogs through grid searches with different learning rates and weight decays. As shown in Table 9, our method outperforms the baseline overall across the datasets, which shows superior transferability and tuning robustness.
| Case | Fine-tuning () |
| 84.5 | |
| 84.6 | |
| Smoothed | 84.4 |
| Cosine distance (Cos) | 84.8 |
| Case | Fine-tuning () |
| 84.7 | |
| 84.8 | |
| 84.7 |
| Method | Dynamic | FT () | |
| Baseline | - | - | 84.3 |
| DTM | ✓ | - | 84.0 (-0.3) |
| DTM | ✓ | ✓ | 85.4 (+1.1) |
5.5 Ablation study
We conduct ablation studies on loss functions, the range of token morphing steps, and target normalization. We also study the impacts of the number of morphing schedules. We pre-train ViT-B/16 with a resolution of for 100 epochs and fine-tune for 100 epochs on ImageNet-1K [39].
5.5.1 Loss function.
We compare various options for the loss function in our method. We compared , , smoothed , and cosine distance. As shown in Table LABEL:tab:ablation_loss_ftn, the model pre-trained using cosine distance outperforms the models with other distance functions.
5.5.2 Range of token morphing iterations.
We compare fine-tuning accuracies of the pre-trained models while varying the ranges that randomly samples the number of morphing iterations in Table LABEL:tab:ablation_step. While our DTM works for all the sampling ranges, K=14 works best.
5.6 Analysis
5.6.1 Ablation study on the importance of dynamics.
Table 11 presents an ablation study on the effectiveness of the dynamic mechanism on DTM. We employ ViT-B/16 [15] with a resolution of . The models are pre-trained for 300 epochs and fine-tuned for 100 epochs on ImageNet-1K [39]. For token morphing without the dynamic mechanism, half of the total 196 tokens are aggregated for each image. As shown in Table 11, token morphing with dynamic mechanism significantly improves the baseline while performance degrades without the dynamic mechanism, demonstrating the importance of dynamic mechanism in DTM.
5.6.2 Faster convergence of DTM.
We analyze the behaviors of models pre-trained by DTM and token-wise objectives in Fig. 5. All the approaches employ CLIP representations for the target. As shown in Fig. 5(a), the fine-tuning accuracies of DTM surpass the accuracy of the token-wise supervision approaches. Fig. 5(b) shows that the model pre-trained by DTM exhibits a lower fine-tuning loss than the baseline model, verifying better convergence. Finally, Fig. 5(c) verifies the consistent effect of DTM across various pre-training epochs.
6 Conclusion
Appendix
This appendix includes additional experimental analyses of our proposed method. First, we present additional examples of spatial inconsistency among patches from pre-trained models in §0.A. Then, we validate the applicability of superpixel algorithms [2] into DTM in §0.B. We also conduct ablation studies on number of morphing schedules and effects of randomness in the number of morphing tokens, randomness in gradual token morphing, and target normalization in §0.C. Finally, §0.D provides implementation details for both pre-training and fine-tuning on ImageNet-1K [39], as well as fine-tuning on ADE20K [55].
Appendix 0.A More Examples on Spatial Inconsistency
We further investigate the spatial inconsistency of patch representations generated by a pre-trained model across various samples. We visualize token morphing results and correct maps of the patch-wise prediction results with and without token morphing using CLIP [35]. Specifically, the correct maps provide information about both correct and incorrect predictions. Patches with correct predictions are highlighted in yellow, while incorrectly predicted patches exhibit results with monotone colors. As shown in Fig. A, patch representations without token morphing reveal spatially inconsistent patch-wise predictions compared to the prediction results with token morphing, which reveals the spatially inconsistent representations among patches. In addition, predictions with token morphing exhibit significantly more correctly predicted patches than predictions without token morphing.
Appendix 0.B Compatibility with Superpixel Algorithms
We perform additional experiments to use the concept of superpixels [2, 8] in ours, both directly to tokens (layer-wise superpixel) and can be combined with DTM. The layer-wise superpixel method uses constant numbers of superpixels and iterations. Table A shows superpixel-based methods enjoy notable gains but do not exceed the bipartite matching one. While the layer-wise superpixel method also utilizes superpixels, the layer-wise token aggregation across the encoder layers bipartite matching approach risks harming intermediate representations during encoding
@@ Appendix 0.B Compatibility with Superpixel Algorithms
| Layer-wise superpixel [2] | 86.7 |
| DTM (Superpixel clustering) [8] | 87.1 |
@@
0.B.1 Ablation Study on the Number of Schedules
Our method can further enhance the diversity of target morphed tokens by employing multiple schedules within a single iteration. Thus, we study the effects of learning diverse morphed tokens derived from multiple morphing schedules simultaneously. We compare the fine-tuning accuracy of models pre-trained by our DTM, varying the number of schedules. We pre-train ViT-B/16 with a resolution of for 100 epochs and fine-tune for 100 epochs on ImageNet-1K [39] and ImageNet-100 [39]. Table B reveals that exploring diverse target morphed tokens improves the representation capability of pre-trained models, leading to increased fine-tuning accuracies of at least 0.1p and 0.2p on ImageNet-1K and ImageNet-100 compared to the model pre-trained by the single DTM approach, respectively. However, the goal of experiencing diverse morphed tokens at once appears to be attained with double scheduling, resulting in no additional gains in performance through further exploration. Given that utilizing two schedules yields the best and is most efficient among all other multiple scheduling options, we adopt the double morphing scheduling approach.
Appendix 0.C Additional ablation studies
0.C.1 Ablation Study on Randomness of DTM.
0.C.1.1 Randomness in the number of morphing tokens.
Our DTM randomizes the number of morphing tokens since fixing the number of morphing tokens does not adequately generate diverse morphed tokens. To verify the effectiveness of the randomness, we compare the fine-tuning accuracies of models pre-trained using DTM with random and fixed numbers of morphing tokens. We pre-train and fine-tune the models for 100 epochs on ImageNet-100 [39]. As reported in the 2nd and 5th rows of Table C, exploring diverse morphed tokens improves the fine-tuning accuracy by 0.5p. This result demonstrates the impact of varying numbers of morphing tokens.
0.C.1.2 Randomness in gradual token morphing.
As morphed tokens can vary by the number of morphing iterations, we apply randomness in token morphing iterations. We compare fine-tuning accuracies of the pre-trained models with random and fixed iteration numbers for morphing to validate the effect of randomness. As shown in the 3rd and 5th rows of Table C, randomly selecting the iteration number for morphing enhances the performance, confirming its effectiveness.
0.C.2 Ablation Study on target normalization.
Target normalization is proven to have a significant impact on MAE [19]. Thus, we verify the impact of target normalization on our DTM. We employ ViT-B/16 with a resolution of for comparison. As shown in the 4th and 5th rows of Table C, target normalization does not yield a positive effect on our DTM, resulting in a decrease in fine-tuning accuracy from 87.9 to 87.7.
Appendix 0.D Implementation details
0.D.0.1 Pre-training on ImageNet-1K.
The pre-training recipe for DTM mainly follows the recipe of BEiT v2 [34]. Table D reports the implementation details for pre-training. We adopt the hyperparameters for ViT-B/16 and ViT-L/16 pre-training from BEiT v2. Specifically, we use layer scales of 0.1 and for ViT-B/16 and ViT-L/16, respectively. We employ both relative positional embeddings and shared relative positional embeddings. The maximum gradient value is constrained to 3.0. We apply color jittering followed by random resizing and cropping for data augmentation. The hyperparameters for ViT-S/16 replicate the settings of ViT-B/16.
0.D.0.2 Fine-tuning on ImageNet-1K.
The fine-tuning recipe for DTM mainly follows the standard supervised learning recipe for ViT. Specifically, we adopt commonly used values for RandAugment, Mixup, Cutmix, and Label Smoothing. On the other hand, we employ relative positional embeddings. Stochastic depth is applied with values of 0.1, 0.1, and 0.2 for ViT-S/16, ViT-B/16, and ViT-L/16, respectively. The overall recipe is detailed in Table E.
@@ Appendix 0.D Implementation details
0.D.0.3 Fine-tuning on ADE20K.
Table F summarizes the fine-tuning recipe of ViT/16 for the semantic segmentation task on ADE20K [55]. We employ AdamW with momentum (0.9, 0.999) and warm-up for 1500 iterations. The learning rate is linearly scheduled with a value of We apply layer-wise learning rate decay of 0.75, stochastic depth of 0.1, and weight decay of 0.05. The model is fine-tuned using 8 V100-32GB GPUs.
0.D.0.4 Transfer learning.
We follow the fine-tuning recipes for DTM to conduct transfer learning to iNaturalist datasets, including iNaturalist 2018, iNaturalist 2019, and mini iNaturalist 2021 [45] and FGVC datasets, including Birds, CUB-200, CIFAR-10, CIFAR-100, and Dogs. However, we additionally perform grid searches of learning rates and weight decay. Specifically, we fine-tune the models with learning rates of , , and and weight decays of 0.05 and 0.1.

### v3

@@ C.1 Ablation Study on the Number of Schedules
| Layer-wise superpixel (Achanta et al., 2012) | 86.7 |
| DTM (Superpixel clustering) (Chang et al., 2023) | 87.1 |
@@
| Case | Fine-tuning () |
| 84.5 | |
| 84.6 | |
| Smoothed | 84.4 |
| Cosine distance (Cos) | 84.8 |
| Case | Fine-tuning () |
| 84.7 | |
| 84.8 | |
| 84.7 |
Appendix D Additional ablation studies
D.1 Ablation Study on Randomness of DTM.
Randomness in the number of morphing tokens. Our DTM randomizes the number of morphing tokens since fixing the number of morphing tokens does not adequately generate diverse morphed tokens. To verify the effectiveness of the randomness, we compare the fine-tuning accuracies of models pre-trained using DTM with random and fixed numbers of morphing tokens. We pre-train and fine-tune the models for 100 epochs on ImageNet-100 (Russakovsky et al., 2015). As reported in the 2nd and 5th rows of Table E, exploring diverse morphed tokens improves the fine-tuning accuracy by 0.5p. This result demonstrates the impact of varying numbers of morphing tokens.
Randomness in gradual token morphing. As morphed tokens can vary by the number of morphing iterations, we apply randomness in token morphing iterations. We compare fine-tuning accuracies of the pre-trained models with random and fixed iteration numbers for morphing to validate the effect of randomness. As shown in the 3rd and 5th rows of Table E, randomly selecting the iteration number for morphing enhances the performance, confirming its effectiveness.
D.2 Ablation Study on target normalization.
Target normalization is proven to have a significant impact on MAE (He et al., 2022). Thus, we verify the impact of target normalization on our DTM. We employ ViT-B/16 with a resolution of for comparison. As shown in the 4th and 5th rows of Table E, target normalization does not yield a positive effect on our DTM, resulting in a decrease in fine-tuning accuracy from 87.9 to 87.7.
D.3 Ablation study
We conduct ablation studies on loss functions, the range of token morphing steps. We also study the impacts of the number of morphing schedules. We pre-train ViT-B/16 with a resolution of for 100 epochs and fine-tune for 100 epochs on ImageNet-1K (Russakovsky et al., 2015).
Loss function. We compare various options for the loss function in our method. We compared , , smoothed , and cosine distance. As shown in Table LABEL:tab:ablation_loss_ftn, the model pre-trained using cosine distance outperforms the models with other distance functions.
Range of token morphing iterations. We compare fine-tuning accuracies of the pre-trained models while varying the ranges that randomly samples the number of morphing iterations in Table LABEL:tab:ablation_step. While our DTM works for all the sampling ranges, K=14 works best.
Appendix E Implementation details
Pre-training on ImageNet-1K. The pre-training recipe for DTM mainly follows the recipe of BEiT v2 (Peng et al., 2022). Table G reports the implementation details for pre-training. We train our framework with ViT-S/16, ViT-B/16, and ViT-L/16 for 300 epochs using AdamW with momentum (0.9, 0.98) and a batch size of 1024. We use a learning rate of with cosine decay and warmup 10 epochs. We employ the CLIP base models (Radford et al., 2021) with its visual projector as a target model across all scales of ViT. Block-wise masking is used with a ratio of 0.4 following (Bao et al., 2021; Peng et al., 2022). Cosine distance is used as a distance metric for the objective according to an ablation study in Appendix. We adopt the hyperparameters for ViT-B/16 and ViT-L/16 pre-training from BEiT v2. Specifically, we use layer scales of 0.1 and for ViT-B/16 and ViT-L/16, respectively. We employ both relative positional embeddings and shared relative positional embeddings. The maximum gradient value is constrained to 3.0. We apply color jittering followed by random resizing and cropping for data augmentation. The hyperparameters for ViT-S/16 replicate the settings of ViT-B/16. For the hyperparameters of DTM, we used 2 for , 1 for , 147 for , and 14 for all . We also pre-train various SSL frameworks through our DTM with the same fundamental setups.
Fine-tuning on ImageNet-1K. We fine-tune our pre-trained models on ImageNet-1K (Russakovsky et al., 2015) by default following the standard protocol (He et al., 2022; Peng et al., 2022). Specifically, pre-trained ViT-S/-B/-L are fine-tuned for 300, 100, and 50 epochs, respectively. Optimization is performed with AdamW using a weight decay of 0.05. We use a layer-wise learning rate decay of 0.6 for ViT-S and ViT-B and 0.8 for ViT-L. Learning rate is set to with a linear warmup for 10 epochs for ViT-S and ViT-B and 5 epochs for ViT-L. We adopt commonly used values for RandAugment, Mixup, Cutmix, and Label Smoothing. On the other hand, we employ relative positional embeddings. Stochastic depth is applied with values of 0.1, 0.1, and 0.2 for ViT-S/16, ViT-B/16, and ViT-L/16, respectively. The overall recipe is detailed in Table H.
@@ Appendix E Implementation details
Fine-tuning on ADE20K. Table I summarizes the fine-tuning recipe of ViT/16 for the semantic segmentation task on ADE20K (Zhou et al., 2017). We employ AdamW with momentum (0.9, 0.999) and warm-up for 1500 iterations. The learning rate is linearly scheduled with a value of We apply layer-wise learning rate decay of 0.75, stochastic depth of 0.1, and weight decay of 0.05. The model is fine-tuned using 8 V100-32GB GPUs.
Transfer learning. We follow the fine-tuning recipes for DTM to conduct transfer learning to iNaturalist datasets, including iNaturalist 2018, iNaturalist 2019, and mini iNaturalist 2021 Van Horn et al. (2018) and FGVC datasets, including Birds, CUB-200, CIFAR-10, CIFAR-100, and Dogs. However, we additionally perform grid searches of learning rates and weight decay. Specifically, we fine-tune the models with learning rates of , , and and weight decays of 0.05 and 0.1.
Applicability on various SSL frameworks. When pre-training and fine-tuning models with MAE (He et al., 2022), BEiT v2 (Peng et al., 2022), and BYOL (Grill et al., 2020), we follow their vanilla training recipes. We pre-train and fine-tune ViT-B/16 for 100 epochs. However, we use CLIP (Radford et al., 2021) target features instead of patchified images to pre-train MAE.